A 33 kg woman climbs up a flight of stairs so she is now 20 meters above the lowest floor. She does this in 25 seconds. How much power does she use?


1
Expert's answer
2021-02-22T10:23:55-0500

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{33\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot20\ m}{25\ s}=258.72\ W."